OpenStudy (anonymous):

what the range for the function f(x)=2(x-4)²+ 3

OpenStudy (haseeb96):

the set of all positive real number is the range of yur question like this x=0 F(0)=2(0-4)^2+3 =2(-4)^2+3 =2(16)+3 =32+3 =35 35 is belong to positive real numbers x=1 F(1)=2(1-4)^2+3 =2(-3)^2+3 =2(9)+3 =18+3 =21 21 is also belong to the positive real number and when we put neagtive number it also become positive For instance x=-3 f(-3)=2(-3-4)^2+3 =2(-7)^2+3 =2(49)+3 =98+3 = 101 101 it is also belong to the positive real no#

OpenStudy (haseeb96):

i think these points clear yur hurdles in this question